Circular No. 6578

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 970228 S. M. Matz, Northwestern University; K. McNaron-Brown, George Ma-

son University; and J. E. Grove and G. H. Share, Naval Research Laboratory, report on behalf of the OSSE team: “The OSSE instrument on the Compton Gamma Ray Observatory had good exposure to the position of GRB 970228 (IAUC 6572) before and after the burst. We are able to derive limits on flux from the source, starting ∼ 33 min after the event, including the period during which BeppoSAX detected an x-ray source (IAUC 6576). No transient or persistent flux has been detected from ∼ 40 keV to 10 MeV. Preliminary post-burst limits (3σ, in units of 10^{−5} photons cm^{−2} s^{−1} keV^{−1}) at ∼ 40–80 keV are as follows: Feb. 28.149–28.444 UT, 5.3; Feb. 28.473–28.630 (roughly the SAX interval), 8.3; and Feb. 28.149–Mar. 3.000, 2.2. The limit on the flux in the day prior to the burst (Feb. 27.124– 28.107) is 3.7 × 10^{−5}. The source position was not continuously observed during these intervals. The OSSE 40–80-keV upper limit during the SAX accumulation is comparable to the flux seen by SAX at 2–10 keV. A single power law, consistent with both the SAX observation and the OSSE limits up to 250 keV, must be softer than photon spectral index −0.3.”

K. Hurley, Space Sciences Laboratory, Berkeley, on behalf of the Ulysses Gamma-Ray Burst Team; E. Costa and M. Feroci, Istituto Astrofisica Spaziale, CNR, Frascati; F. Frontera, Universitá di Ferrara and ITESRE, Bologna; and D. Dal Fiume and M. Orlandini, ITESRE, Bologna, report: “A triangulation of GRB 970228 has been done using preliminary Ulysses and SAX GRBM data. The resulting position is an annulus whose halfwidth is conservatively estimated to be 0^{o}.015 at 3σ confidence, centered at α = 11^{h}03^{m}47^{s}, δ = +36^{o}31′.5 (equinox 2000.0), with central radius 83^{o}.422. Combining this with the BeppoSAX error circle by Costa et al. (IAUC 6572) gives an approximately 6-arcmin^{2} trapezoidal error box with corners at α, δ = 5^{h}01^{m}49^{s}, +11^{o}44′.2; 5^{h}01^{m}46^{s}, +11^{o}45′.2; 5^{h}02^{m}00^{s}, +11^{o}43′.6; 5^{h}01^{m}46^{s}, +11^{o}48′.1. Although this trapezoid formally excludes the fading VLA source reported by Frail et al. (IAUC 6576), the annulus alone is consistent with it and with the MECS/LECS x-ray source reported by Costa et al. Further processing may result in a narrower annulus.”

COMET C/1995 O1 (HALE-BOPP) Naked-eye m_{1} estimates by G. W. Kronk, Troy, IL: Mar. 4.47 UT, 0.3;

6.47, 0.1; 7.47, 0.0; 8.46, 0.0.

1997 March 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green